OpenStudy (anonymous):

wat is 5x squared +100=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Well, if \[5x^2 + 100 = 0\]Then \[5x^2 = -100\]\[\implies x^2 = -20 \implies x = \sqrt{-20} = 2\sqrt{5}i\]

OpenStudy (anonymous):

It s an equation with imaginary roots \[\pm2\sqrt{5}i\]

OpenStudy (anonymous):

\[{5x^2}\] + \[100\] = 0 \[{5x^2}\] = \[-100\] which is equal to \[{x^2}\] =\[\sqrt{20}\] = \[2\] \[\sqrt{5}\]

OpenStudy (anonymous):

by the way Its \[2\sqrt{5i}\]